OpenStudy (anonymous):

COMPLEX VARIABLES: Can someone explain how to find the Arg z and what does it mean?

OpenStudy (anonymous):

If I give you a complex number in \[a+bi\]form, do you know how to convert it to its "polar" form\[re^{i\theta}\]?

OpenStudy (anonymous):

Because when written in its polar form:\[\arg(re^{i\theta})=\theta\]

OpenStudy (anonymous):

where does a and b go?

OpenStudy (anonymous):

I don't know how to convert it to its polar form, could you help me understand that formula?

OpenStudy (anonymous):



OpenStudy (anonymous):

r is the length of the complex number, given by using pythagorean theorem:\[r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

the angle is the argument. You use inverse tangent to get it.\[\theta=\tan ^{-1}(\frac{a}{b})\]

OpenStudy (anonymous):

oh so r = modulus of the point (a,b)?

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

so when asked arg z, it is the theta that they are asking for?

OpenStudy (anonymous):

yes, they want the angle. Be careful when using that formula though. You need to make sure the answer you get reflects the right quadrant. Ex. arg(1+i) vs. arg(-1-i)

OpenStudy (anonymous):

so when asked the arg of 1st quadrant, theta is by the formula you gave me; when asked the arg of quadrant 3, theta is tan inverse of (b/a) ?

OpenStudy (anonymous):

ack! thanks for pointing out my mistake, it should be b/a in the formula i posted.

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

the problem arises because inverse tangent always gives an angle between -pi/2 and pi/2, but you want an angle between 0 and 2pi. So if you were asked for the argument of -1-i, the formula would give pi/4 as an answer, but you know you want something in the 3rd quadrant, so you add pi to get the correct answer of 5\pi/4

OpenStudy (anonymous):

oh ok that makes sense! Thanks!! may i ask another question related? How do i put it into the form of Re^(i theta)? actually what is that formula mean? is it just the form of the vector?

OpenStudy (anonymous):

i understand how to get r and theta now, but i was wondering if it is a vector and why is there the 'e'?

OpenStudy (anonymous):

Once you know the modulus and argument of a complex number, you know what r and theta are. Euler's formula says:\[e^{i\theta}=\cos\theta+i\sin\theta\]

OpenStudy (anonymous):

Note that the modulus of e^(itheta)=1 since:\[|\cos\theta+i\sin\theta|=\cos ^2\theta+\sin^2\theta=1\]So the polar form of a complex number\[re^{i\theta}\]it really just saying, "i want a complex number whose length is r, in the direction of theta"

OpenStudy (anonymous):

oh i get it now!! I really appreciate you helping me!! Thank you so much!!